Solve the equation. 2x +5=8 FIE 0 1 O {-1} ° 0 O {-}
Answer #1

2x+5 sol solve the equation We white 3 2 = 2.2(2x+5) equation as 22 4 2 = 8 Base 2(2x+5) = Hn +10 الري لر are dame H Thus And